Find the projection of the vector A=i-2j+k on the vector B=4i-4j+7k
1
Expert's answer
2019-10-17T11:12:02-0400

The scalar projection is "proj_BA=\frac{A*B}{|B|}=\frac{1*4+(-2)*(-4)+1*7}{\sqrt{4^2+(-4)^2+7^2}}=\frac{19}{9}."